Project

General

Profile

Observing Plan for the December 16-17 2013 Engineering Nights

Proposal Id: 2013A-9999
Exposure Scripts: Most of the exposure script for this night are available in ~DECamObserver/ExposureScripts/engineering_20131216
(from the Observer Console GUI select Exposure Control --> Load Exposure Script --> Choose File)

Alistair will be observing together with Kathy Vivas.

Daytime activities

Cage covers fit check (Freddy)

Best done with the telescope flipped to F/8

Check external photodiode logic (Marco, Klaus)

Test interlock, connection to telemetry DB, connection to SISPI and observer alarms/notifications

LN2 lines 1s and 1r (Alistair)

SISPI software updates (Klaus, Ann)

New GUIs, support for external photodiode, guider mode

(Late) Afternoon

Full set of zeros and dome flats
Script: prenight_calibration.json (note that this is a (standard) DES script. From the the Observer Console GUI select Exposure Control --> Load Exposure Script --> Select DECam Script then choose DES/prenight_calibration.json in the dialog window)

Daytime/Nighttime work

New dome encoder testing (Manuel)

Run only with ion pump (Alistair)

TCS Guiding Offsets (Rolando)

Check out a new TCS version that correctly accounts for guider corrections when reporting the telescope position.

Out of Focus Sky Map (Roberto, Aaron)

Measure astigmatism on center CCDs in preparation of now 4maps control system.
Need to specify details:
  • Hexapod/AOS configuration
  • out of focus value
  • sky positions/script with exposure times and filter

Guider CCD astrometry (lower priority)

We need this to determine the actual location of the guider chips relative to the science array. A necessary step before we can consider to use the "auto" mode of the guider. We tried this last month and it appears that there is a problem with the image data. Something could be wrong with the SISPI guider mode. This needs to be fixed before it makes sense to redo this test.

Use a special readout mode to save long-exposure data from the guider CCDs so that we can determine their astrometric solutions relative to the science array to better than 1". There will be 2 components needed for this test:

Calibration

During the daytime, take zeros and dome flats in each filter using the normal dome-flat exposure times, but with SISPI instructed to save the guider readouts.
The guider calibration script is called guider_zeros_flats.json and can be found in the engineering20131216 directory.
This script does not have the standard action commands to turn of vsub and such. Nor does it include the vsub flush exposures.
Before you run it you have to tell the GCS to use a different directory so that we can keep these images.
As we learned in December, this has to be done when the system is idle, i.e. not taking any exposures.
On the architect console select the GCS. Then, in the command box enter:

set image_directory /data_local/images/fits

Once this script is done you need to switch back using the command (again with the GCS selected)
set image_directory /dev/shm/guider

Also when you configure (or restart) the image directory is switched back. If the timing works out Klaus would like to be online for this to make sure it works.

Guider Exposures

During the nighttime, move to any one of the three coordinates given for the linearity test. Then take 3x20 s exposures in each filter, moving 1 arcmin N of original pointing for 2nd exposure and 1 arcmin E for 3rd exposure in each filter. The guider_astrometry.json script in the engineering20131216 directory is set up to do this.
Like in the previous test you need to switch the image directory so that the images are kept. Using the architect console select the GCS and enter:
As we learned in December, this has to be done when the system is idle, i.e. not taking any exposures.

set image_directory /data_local/images/fits

Once this script is done you need to switch back using the command (again with the GCS selected)
set image_directory /dev/shm/guider

From the November engineering nights

A couple of items from last month- please delete or edit as needed

  • Calibrate Stars around BD+17 4708 (before 01:30UT): Lower priority for this observing run.
  • Deep Y band observations, fringe tests - This is considered lower priority for this engineering run
  • Regression Script - Low priority